# Center-freeness of finite-step solvable groups arising from anabelian geometry

### v1

@@ Abstract.
Anabelian geometry suggests that, for suitably geometric objects, their étale fundamental group determines the object up to isomorphism. From a group-theoretic viewpoint, this philosophy requires rigidity properties of the associated étale fundamental groups, which often follow from their center-freeness. In fact, some profinite groups arising from anabelian geometry are center-free. In the present paper, we investigate how such center-freeness behaves when passing to maximal $m$-step solvable quotients for any integer $m\geq 2$. In particular, we show that the maximal $m$-step solvable quotient of the geometric étale fundamental group of a hyperbolic curve over a field of characteristic 0 is center-free. Furthermore, we show that this implies the injectivity statement, i.e., the rigidity property, of the $m$-step solvable Grothendieck conjecture.
@@ Introduction
Let $G$ be a profinite group. We define the derived series of $G$ by
$G^{[0]}\coloneq G,\qquad G^{[m]}\coloneq[G^{[m-1]},G^{[m-1]}]\qquad(m\geq 1).$
For each $m\in\mathbb{Z}_{\geq 0}$, we set $G^{m}\coloneq G/G^{[m]}$, and call it the maximal $m$-step solvable quotient of $G$. In the present paper, we consider the following property:
Property A.
Let $m\in\mathbb{Z}_{\geq 2}$. Then both $G$ and $G^{m}$ are center-free.
If $G$ is metabelian and center-free, then the natural projection $G\to G^{m}$ is an isomorphism for any $m\in\mathbb{Z}_{\geq 2}$, and hence A holds. In general, however, even if $G$ is center-free, $G^{m}$ need not be center-free. In fact, we can easily construct a counterexample as follows:
  • Let $D_{8}=\langle r,s\mid r^{4}=1,\ s^{2}=1,\ srs^{-1}=r^{-1}\rangle$ be the dihedral group of order 88, and define $\phi:D_{8}\to\operatorname{GL}_{2}(\mathbb{F}_{3})$ by
@@
    Then $G\coloneq(C_{3}\times C_{3})\rtimes_{\phi}D_{8}$ is center-free; however, $G^{2}\cong D_{8}$ is not center-free.
The following are known examples arising from anabelian geometry, which satisfy A:
  • Free pro-$\Sigma$ groups: Free pro-$\Sigma$ groups are center-free. Moreover, the maximal 22-step solvable quotients of free pro-$\Sigma$ groups are also center-free (see, for instance, [1, Section 4]). We can generalize this result from the case $m=2$ to any $m\in\mathbb{Z}_{\geq 2}$.
In the present paper, we give a new example of such a profinite group arising from anabelian geometry. Let $k$ be a field of characteristic 0 with algebraic closure $\overline{k}$, and let $X$ be a smooth curve over $k$. Let $\Sigma$ denote a non-empty set of prime numbers. Write $\pi_{1}^{\mathrm{\acute{e}t}}(X)$ for the étale fundamental group of $X$. Moreover, for any $m\in\mathbb{Z}_{\geq 1}$, we define
@@
Then the following exact sequence holds:
@@
If $X$ is hyperbolic, then $\Delta_{X}$ is center-free (see [14, Proposition 1.6]). For any profinite group $G$, we say that $G$ is ab-torsion-free if the abelianization of each open subgroup of $G$ is torsion-free. Additionally, we say that $G$ is ab-faithful if, for each open subgroup $H$ of $G$ and each open normal subgroup NN of $H$, the natural morphism
$H/N\rightarrow\operatorname{Aut}\bigl(N^{\mathrm{ab}}\bigr)$
induced by conjugation is injective. With the above notation, we have the following theorem:
Theorem A (Propositions 2.5 and 2.6).
Let $m\in\mathbb{Z}_{\geq 2}$. Then every ab-torsion-free, ab-faithful profinite group is center-free, and its maximal $m$-step solvable quotient is also center-free. In particular, if $X$ is hyperbolic, then both $\Delta_{X}$ and $\Delta_{X}^{m}$ are center-free.
The following is a direct corollary of this theorem:
Corollary A (Corollary 2.8).
Let $\mathcal{S}$ be a pro-$\Sigma$ surface groups of genus at least 22 and let $m\in\mathbb{Z}_{\geq 2}$. Then both $\mathcal{S}$ and $\mathcal{S}^{m}$ are center-free.
The same statement for free pro-$\Sigma$ groups of rank $\neq 1$ was proved in [15, Section 1.1]. However, the proof of [15, Proposition 1.1.1] contains an error and does not work as written. In Proposition 1.3, we correct the proof of [15, Proposition 1.1.1]; moreover, in Section 1 we give an explicit computation of the centralizer of a free generator. Throughout the present paper, for a profinite group $G$ and a closed subgroup $H\subset G$, we write $\mathrm{C}_{G}(H)$ for the centralizer of $H$ in $G$. We say that $G$ is slim if $\mathrm{C}_{G}(H)=1$ for every open subgroup $H$ of $G$. In particular, slimness implies that $G$ is center-free.
Theorem B (Theorems 1.5 and 1.6).
Let $\mathcal{F}$ be a (possibly infinitely generated) free pro-$\Sigma$ group of rank $r$ with a free generating set $\mathcal{X}$. Let $m\in\mathbb{Z}_{\geq 2}$. Then, for any nonzero integer $n\in\mathbb{Z}$ and any $x\in\mathcal{X}$, one has
$\mathrm{C}_{\mathcal{F}^{m}}(x^{n})=\langle x\rangle.$
In particular, $\mathcal{F}^{m}$ is slim if $r\neq 1$.
Next, we explain an application of A to the $m$-step solvable analogue of Grothendieck’s conjecture. The original Grothendieck conjecture was first proposed in a letter from Grothendieck to G. Faltings [5], and was proved by S. Mochizuki in [8]. Moreover, in [8, Theorem 18.1], S. Mochizuki proved the following “existence” statement for an $m$-step solvable analogue of the Grothendieck conjecture for hyperbolic curves over a sub-$p$-adic field (i.e., a field that embeds as a subfield of a finitely generated extension of $\mathbb{Q}_{p}$):
Assume $\Sigma=\{p\}$ and that $k$ is a sub-$p$-adic field. Let $m\in\mathbb{Z}_{\geq 2}$. Let $X_{1}$ and $X_{2}$ be smooth curves over $k$. Assume that at least one of $X_{1}$ and $X_{2}$ is hyperbolic. Then for any $G_{k}$-isomorphism
@@
$\operatorname{Isom}_{\overline{k}/k}\bigl(\tilde{X}_{1}^{m}/X_{1},\tilde{X}_{2}^{m}/X_{2}\bigr)\rightarrow\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi^{(m)}_{X_{1}},\Pi^{(m)}_{X_{2}}\bigr)$ (0.1)
is surjective, where $\tilde{X}_{i}^{m}\to X_{i}$ is the maximal geometrically $m$-step solvable pro-$\Sigma$ Galois covering of $X_{i}$, and the right-hand set is the image of the natural map
$\operatorname{Isom}_{G_{k}}\bigl(\Pi_{X_{1}}^{(m+3)},\Pi_{X_{2}}^{(m+3)}\bigr)\rightarrow\operatorname{Isom}_{G_{k}}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr).$
Then A induces the injectivity (i.e., “uniqueness”) statement as follows:
Corollary B (Theorem 2.9).
We keep the notation and assumptions as above. Then the natural map 0.1 is bijective.
In Section 1, we show B. In Section 2, we show A, A and B.
@@
For any profinite group $G$, we define the derived series of $G$ by setting $G^{[0]}\coloneq G$ and, for any $m\in\mathbb{Z}_{\geq 1}$,
$G^{[m]}\coloneq[G^{[m-1]},G^{[m-1]}],$
where $\ast$$\ast$] denotes the closed subgroup topologically generated by commutators. Furthermore, for any $m\in\mathbb{Z}_{\geq 0}$, we set
@@
and call it the maximal $m$-step solvable quotient of $G$. For simplicity, we write $G^{\mathrm{ab}}$ instead of $G^{1}$. Then we have the following lemma:
Lemma A (See [16, Lemma 1.1]).
Let $m,n\in\mathbb{Z}_{\geq 0}$. Let $G$ be a profinite group, and let $H$ be an open subgroup of $G^{m+n}$ containing $(G^{m+n})^{[m]}$. Let $\tilde{H}$ be the inverse image of $H$ in $G$ under the natural surjection $G\twoheadrightarrow G^{m+n}$. Then the natural surjection $\tilde{H}^{n}\twoheadrightarrow H^{n}$ is an isomorphism.
@@ 1. Centralizers in free $m$-step solvable groups
In this section, we compute explicitly the centralizer of a free generator of a free $m$-step solvable pro-$\Sigma$ group. A result of this form is stated in [15, Section 1.1]; however, the proof of [15, Proposition 1.1.1] contains an error and does not work as written. In Proposition 1.3 below, we provide a corrected argument. Throughout this section, let $\Sigma$ be a non-empty set of prime numbers.
Before turning to the proof, we record two ingredients.
$1.$ Pro-$\Sigma$ Fox calculus and the Blanchfield–Lyndon sequence
@@
Let us introduce the pro-$\Sigma$ Fox calculus and the Blanchfield–Lyndon sequence. For a pro-$\Sigma$ group $G$, we define its completed group ring by
@@
    $\partial_{i}(1)=0$, where 11 is the unit of $\mathbb{Z}_{\Sigma}[[\mathcal{F}]]$;
@@
For any $i$, such a free differential is uniquely determined; see [6, Appendix]. Moreover, every $\lambda\in\mathbb{Z}_{\Sigma}[[\mathcal{F}]]$ admits an expansion
@@
$\tilde{\iota}:\mathcal{N}\rightarrow\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]^{\oplus r};\qquad\tilde{\iota}(n)\coloneq\bigl(\pi\circ\partial_{i}(n)\bigr)_{1\leq i\leq r}.$
Since $\pi(n)=1$ for each $n\in\mathcal{N}$, we have $\tilde{\iota}(n_{1}n_{2})=\tilde{\iota}(n_{1})+\tilde{\iota}(n_{2})$. Therefore, $\tilde{\iota}$ is a homomorphism and factors through $\mathcal{N}^{\mathrm{ab}}$. Then we write $\iota$ for the induced morphism
$\mathcal{N}^{\mathrm{ab}}\rightarrow\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]^{\oplus r}.$
@@
Proposition 1.1 (The Blanchfield–Lyndon exact sequence, see [6, Theorem A-2]).
Let $\mathcal{F}$ be a free pro-$\Sigma$ group of finite rank $r$ with free generating set $X=\{x_{i}\}_{1\leq i\leq r}$, and let $\mathcal{N}$ be a closed normal subgroup of $\mathcal{F}$. Then there is an exact sequence
$0\rightarrow\mathcal{N}^{\mathrm{ab}}\xrightarrow{\iota}\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]^{\oplus r}\xrightarrow{f}\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]\xrightarrow{s}\mathbb{Z}_{\Sigma}\rightarrow 0$
of $\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]$-modules, where $f$ is given by
@@
1.2. A first computation of a centralizer in a free $\Sigma pro$duct

### v3

@@ Abstract.
Anabelian geometry suggests that, for suitably geometric objects, their étale fundamental groups determine the geometric objects up to isomorphism. From a group-theoretic viewpoint, this philosophy requires rigidity properties, which often follow from the center-freeness of the associated étale fundamental groups. In fact, some profinite groups arising from anabelian geometry are center-free. For any integer $m\geq 2$, we investigate how such center-freeness behaves under passage to the maximal $m$-step solvable quotients. In particular, we show that the maximal $m$-step solvable quotients of the étale and tame fundamental groups of a hyperbolic curve over a separably closed field are torsion-free and center-free. Furthermore, we show that this implies the rigidity property of the $m$-step solvable Grothendieck conjecture.
@@ Introduction
Let $G$ be a profinite group. We define the topological derived series of $G$ by setting
$G^{[0]}\coloneq G,\qquad G^{[m]}\coloneq\overline{\left[G^{[m-1]},G^{[m-1]}\right]}\qquad(m\geq 1).$
We set $G^{m}\coloneq G/G^{[m]}$ and call it the maximal $m$-step solvable quotient of $G$. Consider the following property:
For any $m\in\mathbb{Z}_{\geq 2}$, the quotient $G^{m}$ is center-free.
Known examples of profinite groups that satisfy the property include:
Free pro-$\Sigma$ groups: Free pro-$\Sigma$ groups are center-free. Moreover, the maximal $2$-step solvable quotients of free pro-$\Sigma$ groups are also center-free (see, for instance, [1, Section 4]). We can generalize this result from the case $m=2$ to all $m\in\mathbb{Z}_{\geq 2}$ immediately.
If $G$ is metabelian and center-free, then for any $m\in\mathbb{Z}_{\geq 2}$ the natural projection $G\to G^{m}$ is an isomorphism, and hence $G^{m}$ is center-free. In general, however, even if $G$ is center-free, the quotient $G^{m}$ need not be center-free. In fact, we can easily construct a counterexample as follows:
Let $D_{8}=\langle r,s\mid r^{4}=1,\ s^{2}=1,\ srs^{-1}=r^{-1}\rangle$ be the dihedral group of order $8$, and define $\phi:D_{8}\to\operatorname{GL}_{2}(\mathbb{F}_{3})$ by
@@
Then $G\coloneq(C_{3}\times C_{3})\rtimes_{\phi}D_{8}$ is center-free; however, the quotient $G^{2}\cong D_{8}$ is not center-free.
In this paper, we give a new example of a profinite group that satisfies the property. Let $k$ be a field (of arbitrary characteristic) with separable closure $\overline{k}$, and let $X$ be a smooth curve over $k$. Note that we always assume that smooth curves are geometrically connected. Let $\Sigma$ denote a non-empty set of prime numbers. We write
$\pi_{1}^{\mathrm{\acute{e}t}}(X,\ast)\qquad(\text{resp. }\pi_{1}^{\mathrm{tame}}(X,\ast))$
for the étale fundamental group (resp. tame fundamental group) of $X$, where $\ast:\operatorname{Spec}(\Omega)\to X$ denotes a geometric point of $X$ and $\Omega$ denotes an algebraically closed field. The fundamental group depends on the choice of base point only up to inner automorphisms, and therefore we omit the choice of base point below.
The first main theorem of this paper is the following:
Theorem A (Theorem 2.9).
Assume that $X$ is hyperbolic, that $k=\overline{k}$, and that $\Sigma$ contains a prime number different from the characteristic of $k$. Then, for any $m\in\mathbb{Z}_{\geq 2}$, the maximal $m$-step solvable quotients of $\pi_{1}^{\mathrm{\acute{e}t}}(X)^{\Sigma}$ and $\pi_{1}^{\mathrm{tame}}(X)^{\Sigma}$ are both torsion-free and center-free.
Corollary A (Corollary 2.10).
For any $m\in\mathbb{Z}_{\geq 2}$, the maximal $m$-step solvable quotient of a pro-$\Sigma$ surface group of genus at least $2$ is torsion-free and center-free.
We say that a profinite group $G$ is slim if the centralizer $\mathrm{C}_{G}\left(H\right)$ of each open subgroup $H\subset G$ in $G$ is trivial (see [8, Definition 0.1]). Since slimness is stronger than center-freeness, it is natural to ask whether the center-freeness statement in Theorem A can be strengthened to slimness. At the time of writing, the author does not know whether these groups are slim in general (see Proposition 2.6 for partial results toward slimness). To the best of the author’s knowledge, slimness for the $m$-step solvable quotients is currently known only for free pro-$\Sigma$ groups, as proved in [15, Section 1.1]. However, the argument of [15, Proposition 1.1.1] contains an error and does not go through as written. In Proposition 1.3, we provide a corrected proof of [15, Proposition 1.1.1], and in Section 1 we give a proof of the slimness of the $m$-step solvable quotients of free pro-$\Sigma$ groups as follows:
Theorem B (Theorem 1.5 and Corollary 1.6).
Let $\mathcal{F}$ be a (possibly infinitely generated) free pro-$\Sigma$ group with a free generating set $\mathcal{X}$. Let $m\in\mathbb{Z}_{\geq 2}$. Then, for any nonzero integer $n\in\mathbb{Z}$ and any $x\in\mathcal{X}$, we have
$\mathrm{C}_{\mathcal{F}^{m}}\left(x^{n}\right)=\overline{\left\langle x\right\rangle}.$
In particular, the quotient $\mathcal{F}^{m}$ is slim if $\mathcal{F}\not\cong\mathbb{Z}_{\Sigma}$.
Next, we explain an application of Theorem A to the $m$-step solvable analogue of the Grothendieck conjecture. In the rest of the introduction, we focus only on the case where the field $k$ is a sub-$p$-adic field for some prime number $p$ (i.e., a field that embeds as a subfield of a finitely generated extension of $\mathbb{Q}_{p}$). In particular, the field $k$ has characteristic 0. For simplicity, we write
@@
By construction, we have the following exact sequence:
@@
Here $G_{k}$ denotes the absolute Galois group of $k$.
The original conjecture of A. Grothendieck was first proposed in his letter to G. Faltings [4] and was proved by S. Mochizuki in [7]. Moreover, in [7, Theorem 18.1], S. Mochizuki proved the following “existence” statement for an $m$-step solvable analogue of the Grothendieck conjecture for hyperbolic curves over a sub-$p$-adic field $k$:
Assume $\Sigma=\{p\}$. Let $m\in\mathbb{Z}_{\geq 2}$. Let $X_{1}$ and $X_{2}$ be smooth curves over a sub-$p$-adic field $k$. Assume that at least one of $X_{1}$ and $X_{2}$ is hyperbolic. Then, for any $G_{k}$-isomorphism
@@
$\operatorname{Isom}_{\overline{k}/k}\bigl(\widetilde{X_{1}}^{m}/X_{1},\widetilde{X_{2}}^{m}/X_{2}\bigr)\to\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr)$ (0.1)
is surjective, where $\widetilde{X_{i}}^{m}\to X_{i}$ is the maximal geometrically $m$-step solvable pro-$\Sigma$ Galois covering of $X_{i}$, and the right-hand set is the image of the natural map
$\operatorname{Isom}_{G_{k}}\bigl(\Pi_{X_{1}}^{(m+3)},\Pi_{X_{2}}^{(m+3)}\bigr)\to\operatorname{Isom}_{G_{k}}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr).$
In this paper, we prove the injectivity statement as follows:
Theorem C (Theorem 2.11).
We keep the notation and assumptions as above. Then the natural map (0.1) is bijective.
@@
For any profinite group $G$, we define the topological derived series of $G$ by $G^{[0]}\coloneq G$ and
$G^{[m]}\coloneq\overline{[G^{[m-1]},G^{[m-1]}]}\qquad(m\geq 1),$
where $\overline{[G^{[m-1]},G^{[m-1]}]}$ denotes the closed subgroup topologically generated by commutators of $G^{[m-1]}$. For any $m\in\mathbb{Z}_{\geq 0}$, we set
@@
and call it the maximal $m$-step solvable quotient of $G$. For simplicity, we write $G^{\mathrm{ab}}$ for the abelianization of $G$. With this notation, we have the following basic lemma:
Lemma A.
Let $f\colon G\to Q$ be a morphism of profinite groups. Let $H\subset Q$ be an open subgroup and set $\tilde{H}\coloneq f^{-1}(H)\subset G$. Fix an integer $n\geq 0$. If $\ker(f)\subset\tilde{H}^{[n]}$, then the natural morphism $\tilde{H}^{n}\to H^{n}$ induced by $f$ is an isomorphism of profinite groups.
Proof.
Since profinite groups are compact Hausdorff, the image of a morphism (i.e., continuous homomorphism) is compact, hence closed. In particular, the morphism $f$ sends closed subgroups to closed subgroups. Hence we have $f(\tilde{H}^{[n]})\subseteq H^{[n]}$. Since $f|_{\tilde{H}}\colon\tilde{H}\to H$ is surjective, the restriction $f|_{\tilde{H}^{[n]}}\colon\tilde{H}^{[n]}\to H^{[n]}$ is also surjective. Consider the commutative diagram with exact rows:
[…]
The kernel of the middle vertical morphism $\tilde{H}\to H$ is $\ker(f)$. By assumption $\ker(f)\subseteq\tilde{H}^{[n]}$, the kernel of the left-hand vertical morphism $\tilde{H}^{[n]}\to H^{[n]}$ is also $\ker(f)$. Applying the snake lemma, the right-hand vertical morphism is an isomorphism. ∎
We will frequently apply Lemma A in the setting that $Q\coloneq G^{m+n}$ and $H$ contains $(G^{m+n})^{[m]}$. In this case, Lemma A shows that the natural surjection $\tilde{H}^{n}\twoheadrightarrow H^{n}$ is an isomorphism. This observation is recorded in [16, Lemma 1.1].
@@ 1. Centralizers in free $m$-step solvable groups
In this section, we compute explicitly the centralizer of a free generator in a free $m$-step solvable pro-$\Sigma$ group. A result of this form is stated in [15, Section 1.1]; however, the proof of [15, Proposition 1.1.1] contains an error and does not work as written. In Proposition 1.3 below, we provide a corrected argument. Throughout this section, let $\Sigma$ be a non-empty set of prime numbers. Moreover, for a profinite group $G$ and a subset $S\subset G$, we define
$\mathrm{C}_{G}\left(S\right)\coloneq\{g\in G\mid\forall s\in S,\,gs=sg\}$
and call it the centralizer of $S$ in $G$. (Note that this group is already closed in $G$, and hence profinite.) When $S=\{x\}$, we write $\mathrm{C}_{G}\left(x\right)$ instead of $\mathrm{C}_{G}\left(\{x\}\right)$ for simplicity.
1.1. Pro-$\Sigma$ Fox calculus and the Blanchfield–Lyndon sequence
@@
We recall the pro-$\Sigma$ Fox calculus and the Blanchfield–Lyndon sequence. For a pro-$\Sigma$ group $G$, we define its completed group ring by
@@
    $\partial_{i}(1)=0$, where $1$ is the unit of $\mathbb{Z}_{\Sigma}[[\mathcal{F}]]$;
@@
For each $i$, such a free differential is uniquely determined; see [6, Appendix]. Moreover, every $\lambda\in\mathbb{Z}_{\Sigma}[[\mathcal{F}]]$ admits an expansion
@@
$\tilde{\iota}:\mathcal{N}\to\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]^{\oplus r};\qquad\tilde{\iota}(n)\coloneq\bigl(\pi\circ\partial_{i}(n)\bigr)_{1\leq i\leq r}.$
Since $\pi(n)=1$ for each $n\in\mathcal{N}$, we have $\tilde{\iota}(n_{1}n_{2})=\tilde{\iota}(n_{1})+\tilde{\iota}(n_{2})$. Therefore, the continuous map $\tilde{\iota}$ is a homomorphism and factors through $\mathcal{N}^{\mathrm{ab}}$. We write $\iota$ for the induced morphism
$\mathcal{N}^{\mathrm{ab}}\to\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]^{\oplus r}.$
@@
Proposition 1.1 (The Blanchfield–Lyndon exact sequence; see [6, Theorem A-2]).
Let $\mathcal{F}$ be a free pro-$\Sigma$ group of finite rank $r$ with free generating set $X=\{x_{i}\}_{1\leq i\leq r}$, and let $\mathcal{N}$ be a closed normal subgroup of $\mathcal{F}$. Then the sequence
$0\to\mathcal{N}^{\mathrm{ab}}\xrightarrow{\iota}\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]^{\oplus r}\xrightarrow{f}\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]\xrightarrow{s}\mathbb{Z}_{\Sigma}\to 0$
of $\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]$-modules is exact. Here, the morphism $f$ is given by
@@
1.2. A computation of a centralizer in a free $\Sigma pro$duct